# Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement

### v1

@@ 5.3 Different Components
Phi-3-mini-4k-instruct 79.92 81.51
Mistral-7B-Instruct-v0.2 79.92 81.89
LLaMA3-8B-Instruct 79.92 82.27
LLaMA3-70B-Instruct 79.92 82.80
BioMistral-7B 79.92 81.67
LLaMA3-OpenBioLLM-8B 79.92 82.36
@@
BioClinicalBERT 79.92 82.27
PubMedBERT 80.10 81.87
Med-CPT 78.25 80.87
BioBERT 78.33 80.20
Different ECG Encoders
ResNet18 79.70 81.26
ResNet50 79.92 82.27
ResNet152 80.35 82.98
ViT-Base 76.20 78.12
ViT-Large 76.14 78.04
ViT-Huge 76.31 78.28
@@ 6 Conclusion
In this paper, we introduced FG-CLEP, a novel approach for fine-grained contrastive learning in ECG-text tasks, addressing the critical issue of incomplete ECG reports, particularly the absence of key waveform features. By incorporating LLMs to generate potential waveform features and validating them with our pipeline, we were able to enhance the quality of ECG reports and better capture the diagnostic reasoning process. Additionally, we tackled the challenges associated with multi-label classification and frequent false negatives by introducing a sigmoid-based loss function and a semantic similarity matrix to guide contrastive learning respectively. Experimental results across six ECG datasets, including PTB-XL, CPSC2018, and CSN, demonstrate that FG-CLEP outperforms existing state-of-the-art methods in both zero-shot prediction and linear probing, highlighting its effectiveness in improving ECG classification and facilitating more accurate diagnostic insights.
Limitations
Although our method effectively leverages large language models (LLMs) to recover crucial waveform features from ECG reports under the challenges of hallucinations and the non-bijective relationship between waveform features and diagnoses, there are certain limitations. One of the main drawbacks is the computational cost and time required, especially during the pre-training phase. The volume of ECG data is extensive, and generating accurate waveform features for such large datasets using LLMs is both time-consuming and resource-intensive. Future research should focus on developing more efficient methods for recovering waveform features.
@@ Appendix A Dataset Analysis
MIMIC-ECG The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs from approximately 160,000 unique patients. These diagnostic ECGs utilize 12 leads and are 10 seconds in duration, with a sampling rate of 500 Hz. https://physionet.org/content/mimic-iv-ecg/1.0/.
PTB-XL The PTB-XL ECG dataset is a large dataset of 21,799 clinical 12-lead ECGs from 18,869 patients of 10-second length. There are four subsets with multi-label classification tasks: Superclass (5 categories), Subclass (23 categories), Form (19 categories), and Rhythm (12 categories). Notably, these four subsets have different numbers of samples. https://physionet.org/content/ptb-xl/1.0.3/.
CPSC2018 This publicly accessible dataset comprises 6,877 standard 12-lead ECG records, each sampled at a rate of 500 Hz, with durations ranging from 6 to 60 seconds. The dataset is annotated with 9 distinct labels. http://2018.icbeb.org/Challenge.html.
Chapman-Shaoxing-Ningbo (CSN) This dataset contains 12-lead ECGs of 45,152 patients with a 500 Hz sampling rate. It features multiple common rhythms and additional cardiovascular conditions, all labeled by professional experts. https://physionet.org/content/ecg-arrhythmia/1.0.0/.
@@ Appendix B Pseudo Code
Appendix C Embedding Visualization
We also demonstrate the effectiveness of our representation learning framework by plotting t-SNE visualizations of ECG embeddings produced for PTB-XL ECGs in five classical waveform features. As shown in Fig. 6, our model produces well-clustered representations. Furthermore, as expected, FG-CLEP learns more fine-grained local waveform features of ECGs. Specifically, FG-CLEP clusters ‘prolonged PR interval’ much better than CLEP.
Figure 6: Embeddings visualization of PTB-XL ECGs in 5 waveform features by (a)CLEP and (b)FG-CLEP. Dimension reduced by t-SNE.
Appendix D Running Cases for Generating Fine-Grained Reports
We present three case studies illustrating how the fine-grained reports with waveform features are generated step by step, as shown in Figure 7, 8, 9. The LLM is prompted with two explicit chain-of-thought instructions: ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ followed by ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’
@@ Appendix D Running Cases for Generating Fine-Grained Reports
Appendix E ECG-Text Retrieval
We attempted to use FG-CLEP to retrieve electrocardiograms (ECGs) from the MIMIC-ECG dataset through text. To test our model’s ability to capture fine-grained waveform features, we tested a series of typical waveform features such as ‘RSR’ Pattern,’ ‘Inverted T-waves,’ and ‘Low QRS voltages.’ Figure 10 shows the Top 3 retrieved ECGs with probabilities all greater than 0.99. Our model demonstrated strong capability in retrieving ECGs through waveform feature text, which can lead to two applications: (1) Helping doctors quickly find similar cases or specific ECG patterns, aiding in diagnosis and treatment decision-making; (2) In medical education, text-based retrieval can quickly find typical ECG cases, assisting in teaching and training, thereby improving educational effectiveness.

### v2

@@ 5.4 Semantic Similarity Matrix
LLaMA3-8B-Instruct 80.78 83.03
LLaMA3-70B-Instruct 80.78 83.42
Qwen3-8B 80.78 83.00
Phi-3-mini-4k-instruct 80.78 81.51
Mistral-7B-Instruct-v0.2 80.78 82.12
BioMistral-7B 80.78 82.67
LLaMA3-OpenBioLLM-8B 80.78 83.23
@@
BioClinicalBERT 80.78 83.03
PubMedBERT 80.92 82.07
Med-CPT 78.55 81.21
BioBERT 78.63 81.20
5.5 Different Components
We conducted experiments to evaluate the performance of our framework using different LLMs and text encoders. The results are presented in Table 5. The findings indicate that our framework is robust across various components. Specifically, for different LLMs (Abdin et al., 2024; Jiang et al., 2023a; AI@Meta, 2024; Labrak et al., 2024; Ankit Pal, 2024), larger LLM provide some performance improvements, though the gains are not substantial. While domain-specific LLMs possess more medical knowledge, our method requires formatting the waveform feature outputs, an area where domain-specific models are less effective, resulting in performance that does not surpass general-purpose models. For different text encoders (Alsentzer et al., 2019; Gu et al., 2021; Jin et al., 2023; Lee et al., 2020), our framework consistently achieves significant improvements.
@@ 6 Conclusion
In this paper, we introduced FG-CLEP, a fine-grained ECG-text contrastive learning framework that enhances waveform understanding by aligning specific ECG segments with report tags. Our coarse-to-fine training process leverages large language models to recover missing waveform features and incorporates a semantic similarity matrix to mitigate false negatives. Extensive experiments on six datasets demonstrate that FG-CLEP achieves state-of-the-art performance in both zero-shot and linear probing settings. These results highlight the effectiveness and generalizability of FG-CLEP.
@@ Appendix A Dataset Analysis
MIMIC-ECG The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs from approximately 160,000 unique patients. These diagnostic ECGs utilize 12 leads and are 10 seconds in duration, with a sampling rate of 500 Hz.
PTB-XL The PTB-XL ECG dataset is a large dataset of 21,799 clinical 12-lead ECGs from 18,869 patients of 10-second length. There are four subsets with multi-label classification tasks: Superclass (5 categories), Subclass (23 categories), Form (19 categories), and Rhythm (12 categories). Notably, these four subsets have different numbers of samples.
CPSC2018 This publicly accessible dataset comprises 6,877 standard 12-lead ECG records, each sampled at a rate of 500 Hz, with durations ranging from 6 to 60 seconds. The dataset is annotated with 9 distinct labels.
Chapman-Shaoxing-Ningbo (CSN) This dataset contains 12-lead ECGs of 45,152 patients with a 500 Hz sampling rate. It features multiple common rhythms and additional cardiovascular conditions, all labeled by professional experts.
@@ Appendix B Pseudo Code
Appendix C Running Cases for Generating Fine-Grained Reports
We present three case studies illustrating how the fine-grained reports with waveform features are generated step by step, as shown in Figure 6, 7, 8. The LLM is prompted with two explicit chain-of-thought instructions: ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ followed by ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’
@@ Appendix C Running Cases for Generating Fine-Grained Reports
Appendix D ECG-Text Retrieval
We attempted to use FG-CLEP to retrieve electrocardiograms (ECGs) from the MIMIC-ECG dataset (Gow et al., ) through text. To test our model’s ability to capture fine-grained waveform features, we tested a series of typical waveform features such as ‘RSR’ Pattern, ‘Inverted T-waves,’ and ‘Low QRS voltages.’ Figure 9 shows the Top 3 retrieved ECGs with probabilities all greater than 0.99. Our model demonstrated strong capability in retrieving ECGs through waveform feature text, which can lead to two applications: (1) Helping doctors quickly find similar cases or specific ECG patterns, aiding in diagnosis and treatment decision-making; (2) In medical education, text-based retrieval can quickly find typical ECG cases, assisting in teaching and training, thereby improving educational effectiveness.